OpenStudy (anonymous):

How do i solve a question like: After earning interest, the balance of an account is $420. The new balance is 7 over 6 of the original balance. How much interest was earned Show your working. P.S I will give a medal for best answer :)

OpenStudy (anonymous):

i think most of us are here not for medals

OpenStudy (anonymous):

anyways regarding the answer, let x be the original amount as per the ques tion (7/6) x= 420 thus x=420*6/7=360 thus the interest earned is 420 -360=60

OpenStudy (anonymous):

(7/6)x=420 Find x to find the original balance and minus it from the current balance to find the interest earned.

OpenStudy (anonymous):

alternatively the interst is (1-6/7)*420 =1/7*420=60

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

welcome , but i request not to ask questions satating donation of medals..